Company KSU Inc., has a leading EPS of $3, and a book value of equity of $34,943 million. The median peer group P/E is 18.75. Assuming there are no differences in the fundamentals among the peer group and the company 4M, find an estimate of the company stock price using the price multiples.

Question 22 options:

1)

$44.05

2)

$56.25

3)

$66.35

4)

$71.10
Answer #1

Price of Company KSU inc stock =P/E*EPS =18.75*3 =56.25
(Option 2 is correct option)